OpenStudy (anonymous):

if x^2+y^2=56 and x-y=4,what is the average of x and y?

OpenStudy (anonymous):

plz me solution

myininaya (myininaya):

(x-y)^2=4^2 (x-y)^2=16 x^2-2xy+y^2=16 -2xy+x^2+y^2=16 -2xy+56=16 -2xy=16-56 -2xy=-40 xy=20 y=20/x does this help iman?

myininaya (myininaya):

so we have y=20/x then x-y=4 can be rewritten as x-20/x=4 x^2-20=4x x^2+4x-20=0 we can find x then plug into y=20/x to find y then we can find (x+y)/2

OpenStudy (anonymous):

(x - y)^2 = x^2 + y^2 -2xy 16 = 56 -2xy 2xy = 40 xy =20 you want average x +y ---- right? 2 (x+y)^2 = x^2 + y^2 +2xy (x + y)^2 =96 x + y = sqrt(24) ----- 2

myininaya (myininaya):

nice! :)